solve the following equation for 0<x<360

a) 6cos² x + sin x - 4

b)9 tan x + tan² x = 5 sec² x - 3


1
Expert's answer
2021-01-11T17:32:28-0500

a) "6cos^2x+sinx-4=0"

let's use the following formula

"cos^2x+sin^2x=1"

Hence, "cos^2x=1-sin^2x"

Then, the equation can be written as follows:

"6(1-sin^2x)+sinx-4=0"

Let's assume that "sin x=u" (hence "1\geq u\geq-1" )

Then, the equation can be written as follows:

"6(1-u^2)+u-4=0"

"-6u^2+u+2=0" (it is the quadratic equation).

Let's find its discriminant D:

"D=1^2-4*(-6)*2=1-(-48)=49=7^2>0"

Hence,

"u1=(-1+7)\/(-6*2)=6\/(-12)=-0.5"

"u2=(-1-7)\/(-6*2)=-8\/(-12)=2\/3"

If "sin x = -0.5" then "x1=210\degree" and "x2=330\degree"

If "sinx=2\/3" then "x3=arcsin(2/3)=41.81\degree" and "x4=arcsin(2/3)\degree+180=41.81\degree+180\degree=221.81\degree"

Answer: roots of the equation are following: "210\degree, 330\degree, 41.81\degree, 221.81\degree"

b)

"9tanx+tan^2x=5sec^2x-3"

"9*sinx\/cosx+sin^2x\/cos^2x=5\/cos^2x-3"

Let's multyply both sides of an equation by "cos^2x" :

"9sinx*cosx+sin^2x=5-3*cos^2x"


Let's use the following famous formulas:

"six2x=2sinxcosx" and "cos2x=cos^2x-sin^2x=1-2sin^2x=2cos^2x-1"

Hence,

"9*sin2x\/2=5-3cos^2x-(1-cos^2x)=4-2cos^2x=4-(2cos^2x-1)-1"

Then,

"4.5*sin2x=3-cos2x" This equation can be rewritten as follows:

"(4.5*sin2x)^2=(3-cos2x)^2"

Hence,

"20.25sin^22x=9-6cos2x+cos^22x"

Then,

"20.25-20.25cos^22x=9-6cos2x+cos^22x"

"-21.25cos^22x+6cos2x+11,25=0"

Let's assume that "cos 2x=u" (hence "-1\leq u \leq 1" )

Then, the equation can be written as follows:

"-21.25u^2+6u+11.25=0" (it is the quadratic equation).

Let's find its discriminant D:

"D=36-4*11.25*(-21.25)=992.25=31.5^2>0"

Hence,

"u1=(-6+31.5)\/(2*(-21.25))=-0.6"

"u2=(-6-31.5)\/(2*(-21.25))=0.8824"


if "cos2x=-0.6" then "2x=arccos( -0.6)=323.14\degree" or "216.86\degree"

hence "x = 161.57\degree" or "x=108.43\degree"


if "cos2x=0.8824" then "2x=arccos(0.8824)= 28.06\degree" or "151.94\degree"

hence "x=14.03\degree" or "x=75.97\degree"

The potential roots of the equation are following: "161.57\degree, 108.43\degree, 14.03\degree, 75.97\degree".

In the solution both parts of the equation were squared. hence, It is necessary to check the occurence of extraneous roots.

Checking the root "161.57\degree"

"9(-0.333)+(-0.333)^2=5(-1.054)^2-3" is not true, hence the root "161.57\degree" is not correct.

Checking the root "108.43\degree"

"9(-3)+(-3)^2=5(-3.163)^2-3" is not true, hence the root "161.57\degree" is not correct.

Checking the root "14.03\degree"

"9(0.249)+(0.249)^2=5(1.03)^2-3" is true, hence the root "161.57\degree" is correct.

Checking the root "75.97\degree"

"9(4.001)+(4.001)^2=5(4.125)^2-3" is not true, hence the root "161.57\degree" is not correct.

Answer: the root of equation is "161.57\degree"